OpenStudy (anonymous):

If anyone knows how to solve for centroids, centers of mass, and moments of inertia...lemme know?

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

single or multivariable calculus...

OpenStudy (anonymous):

multivariable i believe. 2 problems. first one, im given 3 coordinates, each one with a value m. solve for My, Mx, and the center of mass.

OpenStudy (anonymous):

second one, im given a shape on a graph and rho. solve for My, Mx, and center of mass

OpenStudy (anonymous):

ok, so what exactly are you having difficulty with?

OpenStudy (anonymous):

im not given any functions f(x) or g(x), so how would i apply the 4 formulas

OpenStudy (anonymous):

so for instance, on the first one, the center of mass, you aren't given rho(x,y) = ???

OpenStudy (anonymous):

ok lemme copy and paste...sec

OpenStudy (anonymous):

The masses mi are located at the points Pi. Find the moments Mx and My and the center of mass of the system. m1 = 4, m2 = 7, m3 = 9; P1 = (1, 3), P2 = (3, −3), P3 = (−2, −2) solve for Mx, My, and find center of mass.

OpenStudy (anonymous):

again, this is easy cake if im given the functions (formulas ftw lol) buttt...im not so :(

OpenStudy (anonymous):

ybar = 4(1) + 7(3) + 9(-2) = -13/20 xbar = 4(3) + 7(-3) +9(-2) = -27/20 total mass = 20

OpenStudy (anonymous):

Mx = numerator of xbar; My = numerator of ybar. center of course is (xbar, ybar).